OpenStudy (s117):

Which number is prime? A. 63 B. 57 C. 59 D. 51

OpenStudy (anonymous):

Prime numbers can only be divisible by 1 and itself.

OpenStudy (s117):

d

OpenStudy (anonymous):

No, 51 is divisible by 3.

ganeshie8 (ganeshie8):

A number is divisible by 3 if the sum of its digits are divisible by 3

OpenStudy (anonymous):

Oh yeah.. :P 51 = 5 + 1 = 6 6 / 3 = 2

OpenStudy (anonymous):

So which one do you think is correct? @S117

OpenStudy (s117):

i think a,

OpenStudy (anonymous):

No..that is also divisible by 3. 6 + 3 = 9 9 / 3 = 3

OpenStudy (anonymous):

59

OpenStudy (anonymous):

Look at B. 57 = 5 + 7 What's 5 + 7?

OpenStudy (s117):

12

OpenStudy (anonymous):

Yes, and is 12 divisible by 3?

OpenStudy (s117):

yeal

OpenStudy (s117):

yea

OpenStudy (anonymous):

Yes, so it is not prime.

OpenStudy (anonymous):

Let's look at 59 What's 5 + 9?

OpenStudy (anonymous):

Its 59

OpenStudy (anonymous):

@shaunte1234 Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. http://openstudy.com/code-of-conduct

OpenStudy (s117):

14

OpenStudy (s117):

@iGreen.

OpenStudy (anonymous):

Yes, and is 14 divisible by 3?

OpenStudy (s117):

no

OpenStudy (anonymous):

Yes, so 69 is your answer.

OpenStudy (anonymous):

Just joined 15 min. ago sorry I didn't relise there was a code of conduct I'll be sure to follow it but, yeah it is 59

OpenStudy (anonymous):

Yes, you were supposed to read it before you joined. You checked the box and agreed to it.

OpenStudy (anonymous):

its 57 because its like even and odd numbers